OpenStudy (anonymous):

If you deposit $1,000 into a bank account that pays 4% interest compounded continuously, how long will it take the account to grow to $2,000?

OpenStudy (goformit100):

use simple interest formula.

OpenStudy (anonymous):

And what would be a simple interest formula?

OpenStudy (anonymous):

You use Pe^(r)(t), don' t you?

OpenStudy (precal):

you can not use simple interest formula since the compounding is continuous

OpenStudy (anonymous):

Pe^(r)(t) is the continuous compounding formula, isn't it?

OpenStudy (precal):

yes it is so use that one

OpenStudy (anonymous):

ok thank you

OpenStudy (el_tucan):

i don't know if this is right but I used A = Pe^rt right, set A to 2000, divide by a thousand, then ln to solve for t, yes, no?

OpenStudy (anonymous):

I used 2000=1000(1.04)^t and then solved for t.

jimthompson5910 (jim_thompson5910):

edmag16, the equation 2000=1000(1.04)^t only works if you are compounding interest annually however, you want to compound interest continuously

OpenStudy (precal):

just solve for t